OpenStudy (anonymous):

for a medal and fan The legs of a right triangle are 5 units and 15 units. What is the length of the hypotenuse? 200 units square root 200 units square root 250 units 250 units

OpenStudy (mathmate):

Pythagoras Theorem!

OpenStudy (anonymous):

a^2+b^2=c@, so sqrt of 200= 15.81

OpenStudy (ericanoel912):

5^2= 25 15^2= 225 250

OpenStudy (anonymous):

250 sqrt tho right?

OpenStudy (ericanoel912):

now you sqrt

OpenStudy (anonymous):

15.8 right

OpenStudy (ericanoel912):

yep, but it says for c sqrt of 250 so u don't have to actually find the sqrt

OpenStudy (anonymous):

so it would just be 250 without the sqrt right

OpenStudy (ericanoel912):

no.. it is but the answer choice for c says sqrt of 250. that's the answer. you don't have to simplify cause its multiple choice

OpenStudy (anonymous):

ok cool thanks so much i appreciate it! :)